# Tactile sensing enables vertical obstacle negotiation for elongate many-legged robots

### v1

@@ Abstract
Many-legged elongated robots show promise for reliable mobility on rugged landscapes. However, most studies on these systems focus on motion planning in the 2D horizontal plane (e.g., translation and rotation) without addressing rapid vertical motion. Despite their success on mild rugged terrains, recent field tests reveal a critical need for 3D behaviors (e.g., climbing or traversing tall obstacles) in real-world application. The challenges of 3D motion planning partially lie in designing sensing and control for a complex high-degree-of-freedom system, typically with over 25 degrees of freedom. To address the first challenge, we propose a tactile antenna system that enables the robot to probe obstacles and gather information about the structure of the environment. Building on this sensory input, we develop a control framework that integrates data from the antenna and foot contact sensors to dynamically adjust the robot’s vertical body undulation for effective climbing. With the addition of simple, low-bandwidth tactile sensors, a robot with high static stability and redundancy exhibits predictable climbing performance in complex environments using a simple feedback controller. Laboratory and outdoor experiments demonstrate the robot’s ability to climb obstacles up to five times its height. Moreover, the robot exhibits robust climbing capabilities on obstacles covered with flowable, robot-sized random items and those characterized by rapidly changing curvatures. These findings demonstrate an alternative solution to perceive the environment and facilitate effective response for legged robots, paving ways towards future highly capable, low-profile many-legged robots.
@@ I Introduction
Figure 1: Robot climbing in different environments. (Top) The robot climbs a large rock, four times its height, in a confined space with terrain covered in mud, grass, and scattered boulders. (Middle) The robot successfully navigates confined environments with vertical obstacles, unstable metal bars, and flowable plastic disks. (Bottom) The robot climbs obstacles five times its height with rapidly changing curvatures in a laboratory setting.
@@
Despite the challenges in many-legged robots, extensive climbing studies have been conducted on other few-legged robots[20, 28, 6, 16, 34, 41, 43]. Quadrupedal robots accomplish impressive parkour tasks by incorporating vision into their control frameworks [20, 28, 6]. Hexapod and quadrupedal robots utilizing adhesion mechanisms, such as vacuum pumps [47, 13, 29], magnetic adhesion [14, 23, 21], and bio-inspired claws [41, 43, 16], have demonstrated robust climbing capabilities on vertical walls and trees.
@@
However, directly transferring the few-legged climbing behaviors to many-legged robots presents substantial challenge. The first challenges is the vision quality. Climbing in quadrupedal robots rely heavily on camera data to reconstruct the geometric features of obstacles for motion planning [39, 20, 6]. The vision in many-legged robots suffers from the inherent poor quality because (i) many-legged robots often operate in confined environments where low-light conditions obstruct vision [50, 40, 52], (ii) many-legged robots experience substantial body orientation oscillation during locomotion, which significantly compromise the focus of the vision, and (iii) the position of camera is typically positioned too close to the ground, limiting the ability to gather whole-terrain information. The second challenge is the adhesion design. Adhesion-based robots are constrained by their specialized foot designs, which limit their deployment to specific surfaces, such as metal pipes [14, 23, 21] or smooth glass walls [47, 13, 29]. To enable many-legged robots to traverse more general terrains, adhesive foot designs are impractical for enhancing climbing capabilities. Finally, 3D motion planning is particularly challenging for these complicated high degree-of-freedom (DoF) systems, which have over 25 DoF. That is, the complexity of their dynamic, whole-body interactions with the environment poses additional difficulties in developing robust control strategies. Partially because of a lack of systematic locomotion research on their biological counterparts, we have limited intuition on “what to sense” and “how to respond” in many-legged systems.
@@
In addition to vision, tactile sensing offers a reliable short-range perception framework [5, 51, 11]. Recent studies [46, 25, 2, 51, 38] highlight its ability to accurately estimate the geometric features of objects at close range. Compared to vision-based approaches, tactile sensing provides advantages such as lower computational requirements and insensitivity to lighting conditions, despite its limited detection range.
@@
Building on these advantages, recent studies show that tactile sensing has been integrated into mid-sized robotic systems for environment geometry estimation and terrain roughness assessment. Bio-inspired tactile antennas [31, 32, 24] have been successfully integrated into robotic control frameworks, enabling behaviors like wall-following. Integrated with hexapod systems [33, 27, 46], tactile sensing facilitates effective gait adaptation for navigating challenging environments. Additionally, mid-sized many-legged robots [17, 19, 18] have incorporated tactile foot contact sensors into their control frameworks for gait adaptation, demonstrating significant improvements in speed when navigating rugged terrain. Therefore, a tactile sensory system is better suited for many-legged robots to perform climbing tasks, especially on highly rugged terrains with low-light condition.
@@
Laboratory and outdoor experiments demonstrate the robot’s ability to climb obstacles up to five times its center height. Moreover, it successfully navigates obstacles with rapidly changing curvatures and those covered with flowable, robot-sized debris. We also validate the robustness of this climbing controller in complex outdoor environments: the robot successfully completed a pipe inspection within a 0.2 m radius pipe filled with robot-sized rocks and leaves and climbed out of a 0.3 m gap under a bridge, maneuvering over scattered giant rocks, dense weeds, and other vegetation.
@@ II Background: Wave templates in many-legged robot
Figure 3: Tactile sensory system. A. Antenna. 1. The base of the antenna is attached to a Force Resistive Sensor (FSR) using screws. The inner part of the antenna is connected to its tip with two springs, allowing it to deform upon contact with obstacles, thereby preventing jamming. 2. Antenna contact states: A value of 0 represents no contact (void state), while 1 indicates an obstacle has been detected. B. Binary limb contact sensing system. 1. Design of a binary contact sensor for each foot, based on capacitive sensing. 2. Contact state of the leg: 0 indicates no contact, while 1 indicates contact.
@@
The gait of the many-legged robot is determined by the phases of contact ($\phi_{c}$) and lateral body undulation ($\tau_{b}$), combining leg and body waves. Effective coordination, ensuring proper leg retraction during motion, is achieved when $\phi_{c}=\tau_{b}-(\xi/N+1/2)\pi$.
Vertical body undulation is introduced through a wave propagated along the robot’s backbone:
@@
III Tactile Sensory system design
@@ III-A Tactile Antenna design
We developed a bio-inspired antenna equipped with two Force Sensing Resistor (FSRs) sensors capable of detecting forces within a range of 0 to 10 N. As shown in Fig. 3.A.1, each FSR is mounted on a flat surface, with the base of the antenna attached to the opposite side of the FSR using screws. The inner part of the antenna is connected to its tip via two springs, allowing it to deform upon contact with obstacles, thereby preventing jamming.
@@ IV-B1 Obstacle height estimation using antenna
As shown in Figure 5, we use the antenna’s “Hit” (1) and “Void” (0) information (Fig. 3.A.2) to estimate the obstacle’s position relative to the head segment joint. An array $H=[h(t-T+1),h(t-T+2),...,h(t)]$ is created to store antenna’s contact history, where $T$ represents the time interval for sampling the antenna’s hit information. For example, at time $t$, if the antenna detects a hit, then $h(t)=1$ is appended to the contact history array $H$. Otherwise, $h(t)=0$ is appended to $H$. This process generates a binary sequence that records the antenna’s hit history over time.
@@
where $L_{1}$ and $L_{a}$ are the lengths of the head segment and the antenna, respectively, and $\theta_{v,1}$ and $\theta_{a}$ are their corresponding joint angles. Since the length ratio between the tip and the base is less than 1:5, the error caused by tip deformation can be safely ignored. Similar to $H$, an array $Z=[z(t-T+1),z(t-T+2),...,z(t)]$ is created to store the Z-position history of the antenna.
@@ IV-B2 Vertical body undulation control
In the raise up phase, the controller raises the head segment above the obstacle (Fig. 6.A1) and pitches it downward to land the legs on the top of obstacle (Fig. 6.A2). Then, in the drag phase, the controller keeps the belly of the head segment closely follow the obstacle’s top so that robot’s leg can make contact with the ground. Making consistent contact with the ground is important, as the robot gains thrust by periodically making contact with the ground [9, 8]. This behavior is achieved using a proportional controller based on the history of $H$ and $Z$ from antenna.
Figure 7: Feedback control framework for climbing. Two parallel controllers independently manage the pitch motion of the head segment and a subsequent segment. Using the Z-positions ($Z_{max}$ and $Z_{min}$) of the obstacle measured by the antenna, the head controller adjusts the head segment to closely follow the contour of the obstacle. Based on the duty factors ($D_{1},...,D_{5}$), the pitch-down controller identifies the appropriate floating segment nearest to the edge and pitches it down to accelerate the transition from floating to being positioned on top of the obstacle.
The controller first compares the absolute values of highest ($Z_{max}$) and lowest ($Z_{min}$) Z-position (Fig. 6.A). If $|Z_{max}|$ is larger, then head segment is raised, with its vertical joint angle controlled by $K_{p}(Z_{max}+a)$. Here, $a$ is the distance between the head joint and the ground when the robot is in its initial state (Fig. 5.A). Otherwise, the head segment pitches downward to keep a constant distance between the head joint and the obstacle’s contour, with the joint angle controlled by $K_{p}(|Z_{min}|-a)$.
@@
where $K_{p,1}$ and $K_{p,2}$ are the proportional gains, $\theta_{0}$ is constant value that controls the head to pitch downward when no hit is detected. The constant value $a$ is chosen to account for the offset, as the local frame (Fig. 5) is set at the head segment joint, located $a$ cm above the ground. This constant varies based on the robot’s dimensions.
@@
To accelerate this transition, a localized pitch-down motion of the vertical joint nearest to the obstacle’s edge is employed. The duty factor ($D$) of each segment is monitored using data from contact sensors. Segments with a duty factor below the threshold are identified as floating. Among these, the floating segment with the smallest index is considered closest to the edge. Finally, the vertical joint immediately ahead of this segment is controlled to pitch down (Fig. 6.B). Note that, the smallest index starts from 2, as there is no vertical joint ahead of the 1st segment. The joint angle for the pitch-down motion is given by:
$\theta_{v,i}=A_{p}sin(4t)+\theta_{p},$ (6)
where $\theta_{v,i}$ is the vertical joint angle of pitch down joint, $A_{p}$ is the amplitude of the periodic pitch motion, and $\theta_{p}$ is a negative constant that dominates the pitch-down motion. In some situations, where $i=1$, both the head controller and the pitch-down controller attempt to control the head simultaneously. To resolve this conflict, we assign higher priority to the pitch-down controller, ensuring that the head’s pitch motion is governed by the pitch-down controller in such cases.
@@ IV-B3 Maximal vertical climbing capacity
where $h_{0}$ is the maximum distance between the head joint and the ground over gait cycle. $h_{0}$ functions as an offset that raises the belly above the ground.
@@ IV-B4 Experiment results
In each trial, the robot was positioned with its antenna 5 cm in front of the obstacle. For the experiments on box-shaped obstacles, the robot was operated for 14 motion cycles (a total of 42 seconds). For the experiments on bundled-cylinder obstacles, the robot was operated for 10 motion cycles (a total of 30 seconds). The total displacement of those experiments are plotted in Fig. 9.B. According to the results, the robot successfully climbs a 20 cm box-shaped obstacle, which is just below the 24 cm theoretical limit derived in the previous section. Additionally, we observe that the climbing limit increases when the robot is tested on the bundled-cylinder obstacle. This is because the obstacle is not entirely vertical; its slope raises $h_{0}$ in Eq.8, thus raises the climbing limit.
@@
Figure 11: Robot traverses a pipe. The robot completed a pipe inspection task in 80 seconds within a cylindrical pipe measuring 0.2 m in radius and 3 m in length. The pipe contained randomly placed robot-sized rocks, requiring the robot’s climbing capability, while leaves and pine straw introduced entanglement challenges, increasing task complexity. An onboard sports camera mounted on the robot captured footage for the inspection.

### v2

@@ Abstract
Many-legged elongated robots show promise for reliable mobility on rugged landscapes. However, most studies on these systems focus on planar motion planning without addressing rapid vertical motion. Despite their success on mild rugged terrains, recent field tests reveal a critical need for 3D behaviors (e.g., climbing or traversing tall obstacles). The challenges of 3D motion planning partially lie in designing sensing and control for a complex high-degree-of-freedom system, typically with over 25 degrees of freedom. To address the first challenge regarding sensing, we propose a tactile antenna system that enables the robot to probe obstacles to gather information about their structure. Building on this sensory input, we develop a control framework that integrates data from the antenna and foot contact sensors to dynamically adjust the robot’s vertical body undulation for effective climbing. With the addition of simple, low-bandwidth tactile sensors, a robot with high static stability and redundancy exhibits predictable climbing performance in complex environments using a simple feedback controller. Laboratory and outdoor experiments demonstrate the robot’s ability to climb obstacles up to five times its height. Moreover, the robot exhibits robust climbing capabilities on obstacles covered with shifting, robot-sized random items and those characterized by rapidly changing curvatures. These findings demonstrate an alternative solution to perceive the environment and facilitate effective response for legged robots, paving ways towards future highly capable, low-profile many-legged robots.
Page: https://juntaohe.github.io/Climbing_RSS2025/
@@ I Introduction
Figure 1: Tacitly sensing multilegged robot climbing in different environments. (Top) The robot climbs a large rock, four times its height, in a confined space with terrain covered in mud, grass, and scattered boulders. (Middle) The robot successfully navigates confined environments with vertical obstacles, unstable metal bars, and flowable plastic disks. (Bottom) The robot climbs obstacles five times its height with rapidly changing curvatures in a laboratory setting.
@@
Despite the challenges in many-legged robots, extensive climbing studies have been conducted on other few-legged robots[23, 32, 8, 19, 38, 47, 49]. Quadrupedal robots accomplish impressive parkour tasks by incorporating vision into their control frameworks [23, 32, 8]. Hexapod and quadrupedal robots utilizing adhesion mechanisms, such as vacuum pumps [53, 16, 33], magnetic adhesion [17, 26, 24], and bio-inspired claws [47, 49, 19, 3, 42, 7], have demonstrated robust climbing capabilities on vertical walls and trees.
@@
However, directly transferring the few-legged climbing behaviors to many-legged robots presents substantial challenge. The first challenge is the vision quality. Climbing in quadrupedal robots relies heavily on camera data to reconstruct the geometric features of obstacles for motion planning [44, 23, 8]. The vision in many-legged robots suffers from the inherent poor quality because (i) many-legged robots often operate in confined environments where low-light conditions obstruct vision [56, 46, 57], (ii) many-legged robots experience substantial body orientation oscillation during locomotion, which significantly compromise the focus of the vision, and (iii) the position of camera is typically positioned too close to the ground, limiting the ability to gather whole-terrain information. The second challenge is the adhesion design. Adhesion-based robots are constrained by their specialized foot designs, which limit their deployment to specific surfaces, such as metal pipes [17, 26, 24] or smooth glass walls [53, 16, 33]. To enable many-legged robots to traverse more general terrains, adhesive foot designs are impractical for enhancing climbing capabilities. Finally, 3D motion planning is particularly challenging for these complicated high degree-of-freedom (DoF) systems, which have over 25 DoF. That is, the complexity of their dynamic, whole-body interactions with the environment poses additional difficulties in developing robust control strategies. Partially because of a lack of systematic locomotion research on their biological counterparts, we have limited intuition on “what to sense” and “how to respond” in many-legged systems.
@@
In addition to vision, tactile sensing offers a reliable short-range perception framework [6, 13]. Recent studies [52, 29, 2, 43] highlight its ability to accurately estimate the geometric features of objects at close range. Compared to vision-based approaches, tactile sensing provides advantages such as lower computational requirements and insensitivity to lighting conditions, despite its limited detection range.
@@
Building on these advantages, recent studies have integrated tactile sensing into mid-sized robotic systems for tasks such as environmental geometry estimation and terrain roughness assessment. Bio-inspired tactile antennas [35, 36, 27, 28] have been successfully incorporated into robotic control frameworks, enabling behaviors like wall-following and climbing. Snake robots have demonstrated impressive mobility on complex terrain using tactile sensing [15, 45]. Similarly, tactile sensing has been integrated into hexapod systems [37, 31, 52] to support effective gait adaptation in challenging environments. Many-legged robots [20, 22, 21] have also employed tactile foot contact sensors to adapt their gaits, showing significant improvements in speed over rugged terrain. These advances suggest that a tactile sensory system is particularly well-suited for many-legged robots performing climbing tasks, especially in highly rugged and low-light environments.
@@
Laboratory and outdoor experiments demonstrate the robot’s ability to climb obstacles up to five times its center height. Moreover, it successfully navigates obstacles with rapidly changing curvatures and those covered with shifting, robot-sized debris. We also validate the robustness of this climbing controller in complex outdoor environments: the robot successfully completed a pipe inspection within a 20 cm radius pipe filled with robot-sized rocks and leaves and climbed out of a 30 cm gap under a bridge, maneuvering over scattered giant rocks, dense weeds, and other vegetation.
@@ II Background: Wave templates in many-legged robot
Figure 3: Tactile sensory system. A. Antenna design. A1. The base of the antenna is attached to a Force Resistive Sensor (FSR) using screws. The inner part of the antenna is connected to its tip with two springs, allowing it to deform upon contact with obstacles, thereby preventing jamming. A2. Antenna contact states: A value of 0 represents no contact (void state), while 1 indicates an obstacle has been detected. B. Binary limb contact sensing system. B1. Design of a binary contact sensor for each foot, based on capacitive sensing. B2. Contact state of the leg: 0 indicates no contact, while 1 indicates contact.
@@
The gait of the many-legged robot is determined by the phases of contact ($\tau_{c}$) and lateral body undulation ($\tau_{b}$), combining leg and body waves. Effective coordination, ensuring proper leg retraction during motion, is achieved when $\tau_{c}=\tau_{b}-(\xi/N+1/2)\pi$.
Vertical body undulation is introduced via a wave propagated along the robot’s backbone:
@@
III Tactile sensory systems
@@ III-A Tactile Antenna design
We developed an antenna equipped with two Force Sensing Resistor (FSRs) sensors capable of detecting forces within a range of 0 to 10 N. As shown in Fig. 3.A.1, each FSR is mounted on a flat surface, with the base of the antenna attached to the opposite side of the FSR using screws. The inner part of the antenna is connected to its tip via two springs, allowing it to deform upon contact with obstacles, thereby preventing jamming. According to our experiments, the addition of the antenna does not limit the robot’s physical capabilities, such as speed or maneuverability.
@@ IV-B1 Obstacle height estimation using antenna
As shown in Figure 5, we use the antenna’s “Hit” (1) and “Void” (0) information (Fig. 3.A.2) to estimate the obstacle’s position relative to the head segment joint. An array $H=[h(t-T),h(t-T+1),...,h(t)]$ is created to store antenna’s contact history, where $T$ represents the time interval for sampling the antenna’s hit information. For example, at time $t$, if the antenna detects a hit, then $h(t)=1$ is appended to the contact history array $H$. Otherwise, $h(t)=0$ is appended to $H$. This process generates a binary sequence that records the antenna’s hit history over time.
@@
where $L_{1}$ and $L_{a}$ are the lengths of the head segment and the antenna, respectively, and $\theta_{v,1}$ and $\theta_{a}$ are their corresponding joint angles. Since the length ratio between the tip and the base is less than 1:5, the error caused by tip deformation can be safely ignored. Similar to $H$, an array $Z=[z(t-T),z(t-T+1),...,z(t)]$ is created to store the Z-position history of the antenna.
@@ IV-B2 Vertical body undulation control
In the raise-up phase, the controller lifts the head segment above the obstacle (Fig. 6.A1) and pitches it downward to place the legs on top of the obstacle (Fig 6.A2). In the drag phase, the controller keeps the underside (belly) of the head segment closely aligned with the obstacle’s surface, allowing the legs to maintain ground contact. Consistent contact with the ground is essential, as the robot generates thrust by periodically engaging its legs with the surface [11, 10]. This behavior is achieved using a proportional controller that leverages the history of $H$ and $Z$ values from the antenna.
Figure 7: Feedback control framework for climbing. Two parallel controllers independently manage the pitch motion of the head segment and a subsequent segment. Using the Z-positions ($Z_{max}$ and $Z_{min}$) of the obstacle measured by the antenna, the head controller adjusts the head segment to closely follow the contour of the obstacle. Based on the duty factors ($D_{1},...,D_{5}$) measured by foot contact sensors, the pitch-down controller identifies the appropriate floating segment nearest to the edge and pitches it down to accelerate the transition from floating to being positioned on top of the obstacle.
The controller first compares the absolute values of highest ($Z_{max}$) and lowest ($Z_{min}$) Z-position (Fig. 6.A). If $|Z_{max}|$ is larger, then head segment is raised, with its vertical joint angle controlled by $K_{p}(Z_{max}+a)$. Here, $a$ is the distance between the head joint and the ground when the robot is in its initial state (Fig. 5.A). Otherwise, the head segment pitches downward to keep a constant distance between the head joint and the obstacle contour, with the joint angle controlled by $K_{p}(|Z_{min}|-a)$.
@@
where $K_{p,1}$ and $K_{p,2}$ are proportional gains, and $\theta_{0}$ is a constant that pitches the head downward when no contact is detected. The constant $a$ accounts for the vertical offset between the head joint and the ground, as the local frame (Fig 5) is defined at the head segment joint, located $a$ cm above the ground. The value of $a$ varies depending on the robot’s dimensions.
@@
To accelerate the transition over the obstacle, a localized pitch-down motion is applied to the vertical joint nearest to the obstacle’s edge. The duty factor ($D$) of each segment is monitored using data from contact sensors, and segments with a duty factor below a threshold are identified as floating. Among these, the floating segment with the smallest index is considered closest to the edge. The vertical joint immediately ahead of this segment is then commanded to pitch downward (Fig. 6.B). Note that the index starts from 2, as there is no vertical joint ahead of the first segment.
The pitch-down joint angle is defined as:
$\theta_{v,i}=A_{p}\sin(4t)+\theta_{p},$ (6)
where $\theta_{v,i}$ is the vertical joint angle, $A_{p}$ is the amplitude of the periodic pitch motion, and $\theta_{p}$ is a negative constant that biases the motion downward.
In cases where $i=1$, both the head controller and the pitch-down controller attempt to control the head simultaneously. To resolve this conflict, we assign higher priority to the pitch-down controller, ensuring that the head’s motion is governed by it in such situations.
@@ IV-B3 Maximal vertical climbing capacity
where $h_{0}$ is the maximum distance between the head joint and the ground over gait cycle. Note that, parameter $h_{0}$ functions as an offset that raises the belly above the ground.
@@ IV-C1 Single obstacle
In each trial, the robot was positioned with its antenna 5 cm in front of the obstacle. For the experiments on box-shaped obstacles, the robot was operated for 14 motion cycles (a total of 42 seconds). For the experiments on bundled-cylinder obstacles, the robot was operated for 10 motion cycles (a total of 30 seconds). The total displacement of those experiments are plotted in Fig. 9.B. According to the results, the robot successfully climbs a 20 cm box-shaped obstacle, which is just below the 24 cm theoretical limit derived in the previous section. Additionally, we observe that the climbing limit increases when the robot is tested on the bundled-cylinder obstacle. This is because the obstacle is not entirely vertical; its slope (around $60^{\circ}$) raises $h_{0}$ in Eq.8, thus raises the climbing limit.
@@
Figure 11: Traversal of two consecutive obstacles A. Snapshots showing the robot traversing two obstacles (0.15 m and 0.13 m in height, spaced 0.5 m apart) within 150 s. B. Forward ($x$) and vertical ($z$) displacement across five trials. Shaded regions represent the standard deviation.
IV-C2 Multiple obstacles in close succession
To evaluate the robustness of the feedback controller in handling multiple obstacles, we tested the robot on two consecutive obstacles (0.15 m and 0.13 m in height), spaced 0.5 m apart (Fig. 11.A), ensuring that the tail remained on the first obstacle as the head ascended the second. Both obstacles were covered with metal beams to increase the complexity of the task. The robot successfully traversed both obstacles in all five trials. Tracking results for forward and vertical displacement are shown in Fig.1.B.
IV-C3 Shifting obstacles
We also tested the robot on a pile of shifting cylinders (Fig. 12). The pile covered a 60 cm × 60 cm area and reached a height of up to 20 cm, consisting of 9 cm diameter cylinders with varying lengths (5 – 15 cm). In all five trials, the robot successfully reached the peak and exited the pile. Fig. 12 shows snapshots from a representative trial in which the obstacles collapsed and shifted during traversal.
Figure 12: Robot navigating unstable obstacles. Snapshots show the robot traversing a pile of unstable, shifting obstacles — 9 cm diameter cylinders of varying lengths (5 – 15 cm), covering a 60 cm × 60 cm area and reaching up to 20 cm in height.
IV-C4 Outdoor test
@@
Figure 13: Robot traverses a pipe. The robot completed an “inspection” task in 80 seconds within a cylindrical pipe measuring 0.2 m in radius and 3 m in length. The pipe contained randomly placed robot-sized rocks, requiring the robot’s climbing capability, while leaves and pine straw introduced entanglement challenges, increasing task complexity. A camera mounted on the robot captured footage during the traversal.